shanta got 215 marks in an examination. If she had got 10 marks more, she would have scored 60% of the total marks. What are the maximum marks for the examination
Found 2 solutions by mananth, Maths68:
Answer by mananth(16949)

let total marks be x
(215+10)/x= 60/100
22500=60x
x=375
total marks =375

Answer by Maths68(1474)

Let
Maximum Marks = x
Marks = 215
Total marks to get 60% = 215+10 = 225

60% of x = 225
60/100 * x = 225
3/5 * x = 2250
3*x=225*5
3*x=1125
3*x/3=1125/3
x=375
Maximum Marks = 375